# Conceptualization, Methodology, Software, Validation, Formal analysis, Data Curation, Investigation, Writing – original draft, Writing - Review & Editing, Visualization

### v1

@@ 6.4 Parameter Sensitivity
(c) The impact of pre training epochs on results.
(d) The impact of pre training learning rate on results.
@@ 6.5 Case Study
For example, the molecular structures of Atropine and Ipratropium are very similar, yet they have significant functional differences. Atropine is primarily used to treat conditions such as blurred vision and arrhythmia caused by various factors, while Ipratropium is mainly used to treat COPD (chronic obstructive pulmonary disease) and other respiratory diseases, administered via nebulization. By incorporating the 3D modality, we successfully distinguish between medications such as R01A and A03B, N07A, S01K, and recommend the correct medication to the user. As shown in Figure 5, we observe that the molecular embeddings generated by the original molecular encoder and the dual-modality encoder in low-dimensional display show a more uniform distribution, better distinguishing the roles and functions of different molecules.
@@ 7 Conclusion
This paper introduces our developed medication recommendation model, BiMoRec. By integrating dual-modality molecular representations, the model addresses the structural confusion problem and significantly enhances the distinction between molecules, achieving notable improvements in molecular recognition. Furthermore, through single interactions between molecular knowledge and patient representations, and by distilling substructures at both the visit level and the overall level using the patient’s latest visit representation, the model identifies truly useful substructures, thereby significantly improving accuracy and safety. We conduct a series of rigorous experiments on publicly available clinical datasets to demonstrate the validity of our approach.

### v2

@@ 6.4 Parameter Sensitivity
(c) The impact of pre-training epochs on results.
(d) The impact of pre-training learning rate on results.
@@ 6.5 Case Study
Figure 6: Detailed case studies on molecular 3D modes and multi enhancement in MIMIC-III
@@
For example, the molecular structures of Atropine and Ipratropium are very similar, yet they have significant functional differences. Atropine is primarily used to treat conditions such as blurred vision and arrhythmia caused by various factors, while Ipratropium is mainly used to treat COPD (chronic obstructive pulmonary disease) and other respiratory diseases, administered via nebulization. As shown in Figure 6, by incorporating the 3D modality, we successfully distinguish between medications such as R01A and A03B, N07A, and recommend the correct medication to the user. As shown in Figure 5, we observe the low-dimensional representation of molecular embeddings generated by the original molecular encoder and the introduced bimodal encoder. The more uniform distribution in the PCA (Principal Component Analysis) [16] plot indicates that significant variability is captured, making it easier to distinguish different molecules. Similarly, the cluster distribution in the t-SNE (t-distributed Stochastic Neighbor Embedding) [26] plot also supports this observation.
@@ 7 Conclusion
This paper introduces our developed medication recommendation model, BiMoRec. By integrating dual-modality molecular representations, the model addresses the structural confusion problem and significantly enhances the distinction between molecules, achieving notable improvements in molecular recognition. Furthermore, through single interactions between molecular knowledge and patient representations, and by enhancing substructures at both the visit level and the overall level using the patient’s latest visit representation, the model identifies truly useful substructures, thereby significantly improving accuracy and safety. We conduct a series of rigorous experiments on publicly available clinical datasets to demonstrate the validity of our approach.